Given the following reaction Ca(s) + 2HCl(aq) → CaCl2(aq) + H2(g) If 0.500 g of Ca reacts with 0.850 g HCl and 1.21 g CaCl2 are produced, what is the theoretical yield, the percent yield and the limiting reactant?

1.38 g, 87.7%, Ca limiting

1.38 g, 93.8%, HCl limiting

1.29 g, 87.7%, Ca limiting

1.29 g, 93.8% HCl limiting

1.38 g, 92.8%, Ca limiting

1.29 g, 92.8%, HCl limiting
Answer #1

Moles of Ca = Given Weight of Ca / Molar Mass of Ca = 0.5/40.08 = 0.0124 moles ,(Molar Mass of Ca is 40.08g/mol)

Moles of HCl = Given Weight of HCl / Molar Mass of HCl = 0.85/36.46 = 0.0233 moles ,(Molar Mass of HCl is 36.46g/mol)

Moles of CaCl2 (produced practically) = Given Weight of CaCl2 / Molar Mass of CaCl2 = 1.21 / 110.98 = 0.0109 moles, (Molar Mass of Ca Cl2 is 110.98 g/mol)

Intial Final Ca(s) + 2HCl + Ca Cly(aq) + Hz (9) 0.0124 0.0233 0.0124-00 0.0233-200 oC 3C • Sf Ga(s) is Limiting Reagent thero

Theoretical Yield of CaCl2 = 1.29g, HCl is the limiting reagent, and precent yield is 93.8%.